# Bilevel subsidy-enabled mobility hub network design with perturbed utility coalitional choice-based assignment

### v1

@@ 3.2 Lower-level traveler-operator coalition choice model using PURC based structure
Eq. (8) depicts the flow assignment decisions for all links $l$ with positive flows for each OD pair $s$. The first term consists of the link disutility from both travelers and operators. The link disutility from the operator side is weighted by $\alpha$ and $q_{s}/\bar{q}$. $\alpha$ is a general coefficient that provides a different weight towards operator disutility compared to travelers. A higher weight leads to higher impact on operator perspective. $q_{s}/\bar{q}$ is an additional weight that considers the impact from different OD pair volumes. The flow assignment variable $x_{l,s}$ is considered as the possibility of choosing link $l$ for OD group $s$. Each OD pair shall be treated equally from a traveler perspective to reflect stochastic user-equilibrium (SUE) behavior from the perspective of the coalitions of travelers and operators. The weight $q_{s}/\bar{q}$ is the lever that place more focus on OD pairs with higher OD volume. For OD pair with higher than average flow, this causes the OD pair $s$ having more impact on final assignment decision for operators. If there are two OD pairs $(0,1)$ and $(0,2)$. $(0,1)$ has 1 unit flow and $(0,2)$ has 2 units. Both ODs would flow through a mobility hub. On the traveler side, each OD is equally treated for flow assignment. However, when it comes to operators, OD pair $(0,2)$ has a higher weight in the resource allocation decision from a system-wise perspective. Therefore, a weight of 0.67 is placed for node cost disutility when considering OD pair $(0,1)$, while the weight towards OD pair $(0,2)$ becomes 1.33.
@@
The second term consists of node specific values except the capacity Lagrangian $\beta_{l,s}$. The first element represents the magnitude of attraction of link $l$ in OD pair $s$, which is measured by the difference between the source node Lagrangian and the sink node Lagrangian. The second element represents the disutility from the node cost for MOD operators. For links that do not belong to the access link set leading to MOD nodes $N_{M}$, the second element is dropped. For the the access links leading to MOD nodes, the node cost disutility from the operator side is counted towards the flow assignment, which is also weighted by $q_{s}/\bar{q}$.
@@
Eq. (8) mimics the structure of the optimal link-flow decision variable with non-zero value shown in Fosgerau et al., (2022). This concludes the proof. ∎
@@
Bell, (1995) proved that the congestion effect can be reflected by the capacity Lagrangian in a capacitated network, which is also shown in Liu et al., (2024). We follow the same logic to show that the congestion effect can also be evaluated by capacity Lagrangians in the PURC based assignment model.
@@
The congestion effect of waiting for service at a capacitated MOD node $i\in N_{M}$ is captured by the corresponding Lagrangian $\lambda_{i}$.
@@
Assume a node $i$ is at capacity ($v_{i}=1$). If there is an additional flow $\delta_{s}$ for OD pair $s$ on link $l$, with $i$ being the sink node, the capacity constraint Eq. (1c) is therefore exceeded by $\frac{\delta_{s}q_{s}}{z_{i}}$. Assume the additional flow switches to link $l^{\prime}$ while the chosen downstream links have identical costs compared to the chosen downstream links sourcing from node $i$. The change of objective value measured by Eq. (1a) is written as Eq. (9).
@@
As shown in Eq. (10), when the right-hand-side of the node capacity constraint Eq. (1f) increases by 1, the objective value of the KKT in Eq. (4) increases by $\pi_{i}$. Therefore, we can rewrite Eq. (9) to Eq. (11).
$\frac{d\Phi_{1}}{\delta_{s}q_{s}/z_{i}}=\frac{z_{i}\bar{q}}{q_{s}}(d_{l}(2x_{l,s}+\delta_{s})-d_{l^{\prime}}(2x_{l^{\prime},s}+\delta_{s}))+\frac{z_{i}\bar{q}}{q_{s}}(d_{l}p_{l}-d_{l^{\prime}}p_{l^{\prime}}+d_{l}c^{t}_{l}-d_{l^{\prime}}c_{l^{\prime}}^{t}))+\alpha z_{i}(d_{l}c^{o}_{l}-d_{l^{\prime}}c^{o}_{l^{\prime}})=\pi_{i}$ (10)
$\frac{\bar{q}}{q_{s}}(d_{l}(2x_{l,s}+\delta_{s})-d_{l^{\prime}}(2x_{l^{\prime},s}+\delta_{s}))+\frac{\bar{q}}{q_{s}}(d_{l}p_{l}-d_{l^{\prime}}p_{l^{\prime}}+d_{l}c^{t}_{l}-d_{l^{\prime}}c_{l^{\prime}}^{t}))+\alpha(d_{l}c^{o}_{l}-d_{l^{\prime}}c^{o}_{l^{\prime}})=\frac{\pi_{i}}{z_{i}}$ (11)
@@
$x_{l,s}-x_{l^{\prime},s}=-(p_{l}-p_{l^{\prime}})-(c^{t}_{l}+\frac{\pi_{i}q_{s}}{z_{i}\bar{q}}-c_{l^{\prime}}^{t})-\alpha(c^{o}_{l}-c^{o}_{l^{\prime}})\frac{q_{s}}{\bar{q}}$ (12)
Eq. (12) shows that the congestion effect caused by the capacitated node can be evaluated by the Lagrangian of the node capacity constraint. When the switched links are identical in length, the congestion effect equals to $\frac{\pi_{i}q_{s}}{z_{i}\bar{q}}$. This concludes the proof. ∎
@@ 3.3 Upper level model formulation
We present the upper-level problem in this subsection. The decision-maker of the upper-level problem is the MH platform, which determines the optimal subsidy strategy for the mobility services. The bilevel structure follows the Stackelberg equilibrium principle, with the MaaS platform as the leader that sets centralized pricing strategies, while travelers and operators are followers who make decisions to minimize their disutilities. We use revenue maximization as the upper-level objective. However, other equity or social welfare focuses objective can be applied for the upper level objective. From a revenue maximization perspective, the mobility hub design problem can be formulated as Eq. (13).
$\max_{\mathbf{p}}\quad$ $\Phi_{0}=\sum_{s}\sum_{l}d_{l}p_{l}\,x_{s,l}\,q_{s}$ (13a)
s.t. $\mathbf{x}=\arg\min_{\mathbf{x}}\;\Phi_{1}(\mathbf{p})$ (13b)
$0<p_{l}\leq\hat{p},\quad\forall l\in A_{M}$ (13c)
The upper-level problem finds the optimal pricing strategy that maximizes the total revenue from the MH platform users while the service flows and capacities are determined by the lower-level problem. The model assumes that $\hat{p}$ reflects the existing price offered by the MOD operator, and $p_{l}$ is the subsidized price at the link level. Other pricing schemes can be considered as well, such as keeping the same subsidy rate for all trips arriving at a particular MH (by indexing price by MH), or having the same subsidy rate per operator $m$ by indexing price by operator instead.
We can also change the objective to reflect maximization of passenger-miles with an explicit term for subsidies. Assuming there is a fixed subsidy package $R$ provided to the platform operator and the service price is fixed as $p^{{}^{\prime}}$, then the upper-level problem would becomes a flow maximization objective that optimally allocate subsidies. Therefore, Eq. (13) can be rewritten as Eq. (14).
$\max_{\mathbf{r}}\quad$ $\Phi_{0}=\sum_{s}\sum_{l}d_{l}\,x_{s,l}\,q_{s}$ (14a)
s.t. $\sum_{l}d_{l}r_{l}x_{l,s}q_{s}=R$ (14b)
$\mathbf{x}=\arg\min_{\mathbf{x}}\;\Phi_{1}(\mathbf{p^{{}^{\prime}}-r})$ (14c)
$r_{l}\geq 0,\quad\forall l\in A_{M}$ (14d)
The objective of Eq. (14) is flow maximizaation. The service price travelers need to pay in the lower-level is therefore $p^{{}^{\prime}}-r$ as shown in Eq. (14c). In this variant, the total amount of subsidy is fixed. It reflects a scenario where the policymaker seeks to evaluate trade-offs between level of subsidy and flow, particularly relevant for public MH platforms.
Aside from the PURC modifications, the bilevel problem presented in these two subsections feature several nontrivial differences from the generic MaaS assignment game model from (Liu and Chow,, 2024) and (Liu et al.,, 2024). These include:
@@
    Whereas the stochastic assignment game in (Liu et al.,, 2024) lacks a subsidy and the deterministic one in (Liu and Chow,, 2024) only subsidizes a system that has no stable outcome otherwise, the proposed model uses subsidy as the primary upper level pricing decisions, assuming that the FT operator and MOD operators already have set pricing. As shown in (Liu et al.,, 2024), the stochastic assignment game is guaranteed to have a stable outcome with a single cost allocation corresponding to the optimal flow because of the endogeneity in the coalition choices. A solution in which optimal subsidies are zero suggests there is no need to setup a MaaS platform with fare bundling as there should be sufficient incentive for travelers and operators to interact for last mile access out-of-platform.
@@
    The FT operator serves the dual role of being an operator (with lower level MH capacity decisions) and also a platform (with control of the upper level subsidy decisions).
  • 5.
    The general MaaS platform assignment game assumes there is no preexisting network, so as costs for operators go up, the optimal solution will have everybody move to out-of-platform alternatives. In the proposed model, there is a preexisting FT service, so if costs of operating MHs are too high, out-of-platform flows imply travelers either going to their destinations on their own, or going to the transit stations on their own or via an unsubsidized access mode.
@@ 3.4 Solution method
The bilevel programming problem is nonconvex due to the dependency between leader and follower. The bilevel problem can be reformulated into a single-level constrained optimization problem by incorporating the KKT conditions of Eq. (1) as additional constraints into Eq. (13) or Eq. (14). We use Eq. (13) as illustration and the reformulated single-level problem is written as Eq. (15).
$\max_{\mathbf{p,x,v,\mu,\lambda,\beta,\pi}}\quad$ $\Phi^{\prime}_{0}=\sum_{s}\sum_{l}d_{l}p_{l}\,x_{s,l}\,q_{s}$ (15a)
s.t. $\sum_{l}d_{l}r_{l}x_{l,s}q_{s}=R$ (15b)
$0<p_{l}\leq\hat{p},\quad\forall l\in A_{M}$ (15c)
$x_{l,s}(2d_{l}x_{l,s}+d_{l}\left(p_{l}+c^{t}_{l}+\alpha\frac{c^{o}_{l}q_{s}}{\bar{q}}\right)$
$+\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in{l^{+},l^{-}}}\lambda_{i}q_{s}+\beta_{l,s})=0\quad\forall l\in A,s\in S$ (15d)
$\alpha\frac{z_{i}c_{i}}{\bar{q}}-\lambda_{i}z_{i}+\pi_{i}=0,\quad\forall i\in N_{M}$ (15e)
$\mu_{i,s}(\sum_{l}a_{i,l}x_{l,s}-f_{i,s})=0,\quad\forall i\in N,s\in S$ (15f)
$\lambda_{i}(\sum_{l\in A^{-}_{0,i}}\sum_{s\in S}x_{l,s}q_{s}-z_{i}v_{i})=0,\quad\forall i\in N$ (15g)
$\beta_{l,s}(x_{l,s}-1)=0,\quad\forall l\in A,s\in S$ (15h)
$\pi_{i}(v_{i}-1),\quad\forall i\in N$ (15i)
$\sum_{l\in A}a_{i,l}x_{l,s}=\begin{cases}-1,\quad\forall i=o,\\ 1,\quad\forall i=d,\\ 0,\quad\forall i\in N\setminus\{o,d\}\end{cases}\quad\forall s=(o,d)\in S,\,\forall i\in N$ (15j)
$\sum_{l\in A^{-}_{0,i}}\sum_{s\in S}x_{l,s}q_{s}\leq z_{i}v_{i}\quad\forall i\in N_{M}$ (15k)
$0\leq x_{l,s}\leq 1\quad\forall l\in A,\,\forall s\in S$ (15l)
$0\leq v_{i}\leq 1\quad\forall i\in N_{M}$ (15m)
$\mathbf{p,x,v,\mu,\lambda,\beta,\pi}\geq 0$ (15n)
@@
Eqs. (15d, 15e) are the first order conditions. Eq. (15d) is an augmented first order condition because of the zero-flow links involved in the final solution as indicated in Fosgerau et al., (2022). Eqs. (15f - 15i) are the complementarity slackness conditions. It has been shown (e.g. Dempe, (2002)) that bilevel problems with linear upper level and convex quadratic lower level programs can be reformulated into a single level mathematical program with complementarity constraints (MPCC) that can be solved to global optimality. Sinha et al., (2017) provide an example algorithm based on branch-and-bound.
@@
However, the added KKT multipliers and complementarity constraints pose great scalability challenges, especially when facing a large network. Therefore, it is impractical to solve for a global optimal solution using branch-and-bound. To address this issue, we adopt the gap function-based method proposed by Marcotte and Zhu, (1996) to solve the bilevel problem effectively while still keeping high solution quality. Instead of directly appending the KKT conditions into upper-level constraints, we move part of the high-dimension constraints into the objective function and formulate them as penalty terms. The new penalty added objective and the adjusted constraints are written as Eq. (16).
@@
Eqs. (15b,15c,15e-15m) (16c)
The penalty term $\Lambda_{l,s}$ is the augmented first order condition Eq. (15d) for each l,s{l,s} pair. The penalty coefficient $\rho$ is a tunable hyperparameter that modifies the magnitude of the penalty. The higher the value of $\rho$, the heavier the penalty when not solved towards the lower-level optima. When $\rho$ is sufficiently high, Eq. (16) is equivalent to minimizing the upper level objective while obtaining the lower level optimum. Therefore, we can also use Eq. (16) to obtain a global optimum of this bilevel problem.
@@
The equality constraint Eq. (15d) can be written as Eq. (17).
@@
When $\rho\geq max(\eta)$, the penalty term $\rho\sum\Lambda$ dominate the incentive to violate Eq. (17b). Under this condition, any optimal solution of the original problem is optimal for the penalized problem. The equality constraint Eq. (15d) is recovered at the optimum even though the feasible set only enforces Eq. (18b). This completes the proof. ∎
To further reduce the computational load, we propose another algorithmic improvement. Since the lower-level problem is very efficient to solve due to its convexity, it is easy to obtain the lower-level optimum when p is fixed. We use an iterative update process to exploit this property and reduce the solution time for Eq. (16). The overall proposed algorithm is summarized in Algorithm 1. The hyperparameters $\rho^{0}$, $\psi^{-}$, and $\psi^{+}$ control the rate of penalty update, where $\psi^{+}>1$ and $\psi^{-}<1$. By starting with a small value $\rho^{0}$, Eq. (16) is easier to solve to optimality. This comes with the cost of violating lower-level KKT condition by a large margin ($\sum\Lambda>>\epsilon$). By gradually increasing $\rho$ with $\psi^{+}$, the KKT violation shrinks, leading to the global optimum. However, when $\rho$ becomes too large, the single-level solution process can become computationally intensive. Therefore, we reduce $\rho$ by $\psi^{-}$ to control the step size. Within each iteration, the previously solved $\mathbf{p}$ (or $\mathbf{r}$ if the decision variable is the subsidy) is used to solve the lower-level problem first. This solution, along with all dual values and their bounds, can warm-start the single-level solution algorithm in a tighter space, leading to a shorter runtime. After meeting the convergence criteria or reaching the maximum number of iterations, we output the final values of $(\mathbf{p}^{*},\mathbf{x}^{*},\mathbf{v}^{*})$.
@@
1:Input: Initial upper-level variable $\mathbf{p}^{0}$, initial penalty parameter $\rho^{0}$, bound relaxation coefficient $\zeta$, penalty parameter update coefficient $\psi^{-},\psi^{+}$, optimality gap target $\tau$, objective tolerance $\epsilon$, time limit $T$, maximum iterations $K$
@@
4:  Step 1: For given $\mathbf{p}^{k}$, solve the lower-level problem Eq. (1) to obtain optimal $\mathbf{x}^{k},\mathbf{v}^{k}$

### v2

@@ 3.2 Lower-level traveler-operator coalition choice assignment model using PURC based structure
Eq. (8) depicts the flow assignment decisions for all links $l$ with positive flows for each OD pair $s$. The first term consists of the link disutility from both travelers and operators. The link disutilities from the two sides are weighted by $\alpha_{1},\alpha_{2}$, and $q_{s}/\bar{q}$. $\alpha_{1}$ and $\alpha_{2}$ are traveler and operator dedicated coefficients that provide different weights towards the overall utility objective. The ratio of $\alpha_{1}/\alpha_{2}$ depicts which side having higher impact towards the flow assignment behavior. $q_{s}/\bar{q}$ is an additional weight that considers the impact from different OD pair volumes. The flow assignment variable $x_{l,s}$ is considered as the possibility of choosing link $l$ for OD group $s$. Each OD pair shall be treated equally from a traveler perspective to reflect stochastic user-equilibrium (SUE) behavior from the perspective of the coalitions of travelers and operators. The weight $q_{s}/\bar{q}$ is the lever that places more focus on OD pairs with higher OD volume.
@@
The second term consists of node and capacity specific values. The first element represents the magnitude of attraction of link $l$ in OD pair $s$, which is measured by the difference between the source and sink node Lagrange multipliers. The second element represents the disutility from the cost of serving a node for MOD operators. For links that do not belong to the access link set leading to MOD nodes $N_{M}$, the second element is dropped. For the the access links leading to MOD nodes, the node cost disutility from the operator side is counted towards the flow assignment, which is also weighted by $q_{s}/\bar{q}$. The remaining three terms are the Lagrange multipliers associated with MOD service node capacity, hub capacity, and link flow constraints.
@@
Eq. (8) mimics the structure of the optimal link-flow decision variable with non-zero value shown in Fosgerau et al. (2022). This concludes the proof. ∎
Representing the differences in the joint behavior of the traveler and operator coalitions, $\alpha_{1}$ and $\alpha_{2}$ would typically be fitted to data in practice to help regulate the dispersion effect observed in the data. As shown in Fosgerau et al. (2022), route dispersion intensifies when the first term in Eq. (1a) holds greater relative weight. If the observed flows have less divergent route choice behavior, the magnitudes of $\alpha_{1}$ and $\alpha_{2}$ can be both raised while maintaining the same ratio of $\alpha_{1}/\alpha_{2}$.
@@
Bell (1995) proved that the congestion effect can be reflected by the capacity Lagrange multiplier in a capacitated network, which is also shown in Liu et al. (2025). We follow the same logic to show that the congestion effect can also be evaluated by capacity Lagrange multipliers in the PURC based assignment model.
@@
The congestion effect of waiting for service at a capacitated MOD access link $l\in A_{i}^{\sim},i\in N_{M}$ is captured by the corresponding Lagrange multiplier $\lambda_{l}$. Similarly, the congestion effect caused by the capacitated hub facility $l\in A_{h}^{\sim},h\in H$ is captured by the corresponding Lagrange multiplier $\varphi_{l}$.
@@
We use the congestion effect at MOD access links as an example. Assume an access link $l$ with sink node $i$ is at capacity ($v_{l}=1$). If there is an additional flow $\delta_{s}$ for OD pair $s$ on link $l$, the capacity constraint Eq. (1c) is therefore exceeded by $\frac{\delta_{s}q_{s}}{z_{l}}$. Assume the additional flow switches to link $l^{\prime}$ while the chosen downstream links have identical costs compared to the chosen downstream links sourcing from node $i$. The change of objective value measured by Eq. (1a) is written as Eq. (9).
@@
As shown in Eq. (10), when the right-hand side of the node capacity constraint Eq. (1f) increases by 1, the objective value of the KKT in Eq. (4) increases by $\pi_{l}$. Therefore, we can rewrite Eq. (9) as Eq. (11).
$\frac{d\Phi_{1}}{\delta_{s}q_{s}/z_{l}}=\frac{z_{l}\bar{q}}{q_{s}}(d_{l}(2x_{l,s}+\delta_{s})-d_{l^{\prime}}(2x_{l^{\prime},s}+\delta_{s}))+\frac{z_{l}\bar{q}}{q_{s}}(d_{l}p_{l}-d_{l^{\prime}}p_{l^{\prime}}+d_{l}c^{t}_{l}-d_{l^{\prime}}c_{l^{\prime}}^{t}))+\alpha z_{l}(d_{l}c^{o}_{l}-d_{l^{\prime}}c^{o}_{l^{\prime}})=\pi_{l}$ (10)
$\frac{\bar{q}}{q_{s}}(d_{l}(2x_{l,s}+\delta_{s})-d_{l^{\prime}}(2x_{l^{\prime},s}+\delta_{s}))+\frac{\bar{q}}{q_{s}}(d_{l}p_{l}-d_{l^{\prime}}p_{l^{\prime}}+d_{l}c^{t}_{l}-d_{l^{\prime}}c_{l^{\prime}}^{t}))+\alpha(d_{l}c^{o}_{l}-d_{l^{\prime}}c^{o}_{l^{\prime}})=\frac{\pi_{l}}{z_{l}}$ (11)
@@
$x_{l,s}-x_{l^{\prime},s}=-(p_{l}-p_{l^{\prime}})-(c^{t}_{l}+\frac{\pi_{l}q_{s}}{z_{l}\bar{q}}-c_{l^{\prime}}^{t})-\alpha(c^{o}_{l}-c^{o}_{l^{\prime}})\frac{q_{s}}{\bar{q}}$ (12)
Eq. (12) shows that the congestion effect caused by the capacitated node can be evaluated by the Lagrange multiplier of the node capacity constraint. When the switched links are identical in length, the congestion effect equals to $\frac{\pi_{l}q_{l}}{z_{l}\bar{q}}$. The same logic can be applied to capacitated hub links. This concludes the proof. ∎
@@ 3.3 Upper level model formulation
We present the upper-level problem in this subsection. The decision-maker of the upper-level problem is the MH platform, which determines the optimal OD-based pricing strategy for accessing the mobility service, the capacity assigned to hub locations, and the subsidy provided to incentivize operator participation. The bilevel structure follows the Stackelberg equilibrium principle, with the MaaS platform as the leader that sets pricing, capacity, and subsidy decisions, while travelers and operators are followers who make decisions to minimize their disutilities. We use revenue maximization as the upper-level objective. However, other equity or social welfare focuses objective can be applied for the upper level objective. From a revenue maximization perspective, the mobility hub design problem can be formulated as Eq. (13).
$\max_{\mathbf{p,r,b}}\quad$ $\Phi_{0}=\sum_{s}\sum_{l\in A_{s}^{+}}p_{s}\,x_{l,s}\,q_{s}-\sum_{l\in A_{H}}\sum_{i\in l^{-}}c_{i}b_{l}-\sum_{m\in M}\sum_{i\in N_{m}}\sum_{l\in A_{i}^{\sim}}\sum_{s\in S}r_{l}x_{l,s}q_{s}$ (13a)
s.t. $\mathbf{x}=\arg\min_{\mathbf{x}}\;\Phi_{1}(\mathbf{p,r,b})$ (13b)
$0\leq p_{s}\leq\hat{p},\quad\forall s\in S$ (13c)
$0\leq b_{l}\leq\hat{b},\quad\forall l\in A_{H}$ (13d)
$0\leq r_{l}\leq\hat{r},\quad\forall l\in A_{i}^{\sim},\forall i\in N_{M}$ (13e)
$\sum_{i\in N_{m}}\sum_{l\in A_{i}^{\sim}}\sum_{s\in S}r_{l}x_{l,s}q_{s}-\sum_{l\in A_{m}}\sum_{s\in S}d_{l}c_{l}^{o}x_{l,s}q_{s}-\sum_{i\in N_{m}}\sum_{l\in A_{i}^{\sim}}z_{l}c_{i}v_{l}\geq 0,\quad\forall m\in M$ (13f)
The upper-level problem finds the optimal pricing strategy, along with the hub capacity and operator subsidy, that maximize the total revenue of the MH platform, while the service flows and MOD route capacities are determined by the lower-level problem. The first and second terms represent the total income from passenger access and the hub operating cost. The third term represents the total subsidy provided to participating MOD operators. The model assumes that $\hat{p},\hat{b}$, and $\hat{r}$ reflect the existing conditions that the market accepts. These parameters could be further segmented for specific traveler and operator groups. We adopt uniform, system-wide values for simplicity in Eq. (13c - 13e). Eq. (13f) requires that the subsidy for each MOD operator cover the operating cost so participation in the MH-oriented platform is profitable.
We can also slightly modify the objective so that subsidy also comes from outside resources (e.g., local authority). Assuming there is a fixed subsidy package $R$ provided to the platform operator, then by directly adding $R$ to Eq. (13a), we can effectively capture the impact of outside subsidy without modifying the constraints.
Aside from the PURC modifications, the bilevel problem presented in these two subsections feature several nontrivial differences from the generic MaaS assignment game model from (Liu and Chow, 2024) and (Liu et al., 2025). These include:
@@
    Whereas the stochastic assignment game in (Liu et al., 2025) lacks a subsidy and the deterministic one in (Liu and Chow, 2024) only subsidizes a system that has no stable outcome otherwise, the proposed model determines both price and subsidy as upper-level decisions with a guaranteed stable outcome. As shown in Liu et al. (2025), the stochastic assignment game is guaranteed to have a stable outcome with a single cost allocation corresponding to the optimal flow because of the endogeneity in the coalition choices. The added subsidy further assures the participating operator’s profitability while strictly following the coalition structure, making the model more applicable in real-world adaptation.
@@
    The general MaaS platform assignment game assumes there is no preexisting network, so as costs for operators go up, the optimal solution will have everybody move to out-of-platform alternatives. In the proposed model, there is a preexisting FT service, so if costs of operating MHs are too high, out-of-platform flows imply travelers either going to their destinations or to the transit stations on their own modes via driving, walking, or pay-as-you-go options.
@@ 3.4 Solution method
The bilevel programming problem is nonconvex due to the dependency between leader and follower. The bilevel problem can be reformulated into a single-level constrained optimization problem by incorporating the KKT conditions of Eq. (1) as additional constraints into Eq. (13). The reformulated single-level problem is written as Eq. (14).
$\max_{\mathbf{p,x,v,\mu,\lambda,\beta,\pi}}\quad$ $\Phi^{\prime}_{0}=\sum_{s}\sum_{l\in A_{s}^{+}}p_{s}\,x_{l,s}\,q_{s}-\sum_{l\in A_{H}}\sum_{i\in l^{-}}c_{i}b_{l}-\sum_{m\in M}\sum_{i\in N_{m}}\sum_{l\in A_{i}^{\sim}}\sum_{s\in S}r_{l}x_{l,s}q_{s}$ (14a)
s.t. $0\leq p_{s}\leq\hat{p},\quad\forall s\in S$ (14b)
$0\leq b_{l}\leq\hat{b},\quad\forall l\in A_{H}$ (14c)
$0\leq r_{l}\leq\hat{r},\quad\forall l\in A_{i}^{\sim},\forall i\in N_{M}$ (14d)
$\sum_{i\in N_{m}}\sum_{l\in A_{i}^{\sim}}\sum_{s\in S}r_{l}x_{l,s}q_{s}-\sum_{l\in A_{m}}\sum_{s\in S}d_{l}c_{l}^{o}x_{l,s}q_{s}-\sum_{i\in N_{m}}\sum_{l\in A_{i}^{\sim}}z_{l}c_{i}v_{l}\geq 0,\quad\forall m\in M$ (14e)
$x_{l,s}(2d_{l}x_{l,s}+d_{l}\left(\alpha_{1}\left(p_{l}+c_{l}\right)+\alpha_{2}\frac{(c^{o}_{l}-r_{l})q_{s}}{\bar{q}}\right)$
$+\sum_{i\in{l^{+},l^{-}}}\mu_{i,s}a_{i,l}+\sum_{i\in l^{+}}\alpha_{2}c_{i}\frac{q_{s}}{\bar{q}}+\sum_{l^{\prime-}\in l^{+}}\varphi_{l^{\prime}}q_{s}+\beta_{l,s})=0\quad\forall l\in A,\forall s\in S$ (14f)
$\alpha_{2}\frac{z_{l}c_{i}}{\bar{q}}-\lambda_{l}z_{l}+\pi_{l}=0,\quad\forall l\in A_{i}^{\sim},\forall i\in N_{M}$ (14g)
$\mu_{i,s}(\sum_{l}a_{i,l}x_{l,s}-f_{i,s})=0,\quad\forall i\in N,\forall s\in S$ (14h)
$\lambda_{l}(\sum_{s\in S}x_{l,s}q_{s}-z_{l}v_{l})=0\quad\forall l\in A_{i}^{\sim},\forall i\in N_{M}$ (14i)
$\varphi_{\tilde{l}}(\sum_{l\in A_{i,h}^{+}}\sum_{s\in S}x_{l,s}q_{s}-b_{\tilde{l}})=0,\quad\forall\tilde{l}\in A_{h}^{\sim},\forall h\in H$ (14j)
$\beta_{l,s}(x_{l,s}-1)=0,\quad\forall l\in A,s\in S$ (14k)
$\pi_{l}(v_{l}-1)=0,\quad\forall l\in A_{i}^{\sim},\forall i\in N_{M}$ (14l)
$\sum_{l\in A}a_{i,l}x_{l,s}=\begin{cases}-1,\quad\forall i=o,\\ 1,\quad\forall i=d,\\ 0,\quad\forall i\in N\setminus\{o,d\}\end{cases}\quad\forall s=(o,d)\in S,\,\forall i\in N$ (14m)
$\sum_{s\in S}x_{l,s}q_{s}\leq z_{l}v_{l}\quad\forall l\in A_{i}^{\sim},\forall i\in N_{M}$ (14n)
$\sum_{l\in A^{+}_{i,h}}\sum_{s\in S}x_{l,s}q_{s}\leq b_{\tilde{l}}\quad\forall\tilde{l}\in A_{h}^{\sim},\forall h\in H$ (14o)
$\mathbf{p,x,v,\mu,\lambda,\beta,\pi}\geq 0$ (14p)
@@
Eqs. (14f, 14g) are the first order conditions. Eq. (14f) is an augmented first order condition because of the zero-flow links involved in the final solution as indicated in Fosgerau et al. (2022). Eqs. (14h - 14l) are the complementarity slackness conditions. It has been shown (e.g. Dempe (2002)) that bilevel problems with linear upper level and convex quadratic lower level programs can be reformulated into a single level mathematical program with complementarity constraints (MPCC) that can be solved to global optimality. Sinha et al. (2017) provide an example algorithm based on branch-and-bound.
@@
However, the added KKT multipliers and complementarity constraints pose great scalability challenges, especially when facing a large network. Therefore, it is impractical to solve for a global optimal solution using branch-and-bound. To address this issue, we adopt the gap function-based method proposed by Marcotte and Zhu (1996) to solve the bilevel problem effectively while still keeping high solution quality. Instead of directly appending the KKT conditions into upper-level constraints, we move part of the high-dimension constraints into the objective function and formulate them as penalty terms. The new penalty added objective and the adjusted constraints are written as Eq. (15).
@@
Eqs. (14b-14d,14g-14p) (15c)
The penalty term $\Lambda_{l,s}$ is the augmented first order condition Eq. (14f) for each l,s{l,s} pair. The penalty coefficient $\rho$ is a tunable hyperparameter that modifies the magnitude of the penalty. The higher the value of $\rho$, the heavier the penalty when not solved towards the lower-level optima. When $\rho$ is sufficiently high, Eq. (15) is equivalent to minimizing the upper level objective while obtaining the lower level optimum. Therefore, we can also solve Eq. (15) to obtain a global optimum of this bilevel problem.
@@
The equality constraint Eq. (14f) can be written as Eq. (16).
@@
When $\rho\geq max(\eta)$, the penalty term $\rho\sum\Lambda$ dominates the incentive to violate Eq. (16b). Under this condition, any optimal solution of the original problem is optimal for the penalized problem. The equality constraint Eq. (14f) is recovered at the optimum even though the feasible set only enforces Eq. (17b). This completes the proof. ∎
To further reduce the computational load, we propose another algorithmic improvement. Since the lower-level problem is very efficient to solve due to its convexity, it is easy to obtain the lower-level optimum when p and r are fixed. We use an iterative update process to exploit this property and reduce the solution time for Eq. (15). The overall proposed algorithm is summarized in Algorithm 1. The hyperparameters $\rho^{0}$, $\psi^{-}$, and $\psi^{+}$ control the rate of penalty update, where $\psi^{+}>1$ and $\psi^{-}<1$. By starting with a small value $\rho^{0}$, Eq. (15) is easier to solve to optimality. This comes with the cost of violating lower-level KKT condition by a large margin ($\sum\Lambda>>\epsilon$). By gradually increasing $\rho$ with $\psi^{+}$, the KKT violation shrinks, leading to the global optimum. However, when $\rho$ becomes too large, the single-level solution process can become computationally intensive. Therefore, we reduce $\rho$ by $\psi^{-}$ to control the step size. Within each iteration, the previously solved $\mathbf{p}$ and $\mathbf{r}$ are used to solve the lower-level problem first. This solution, along with all dual values and their bounds, can warm-start the single-level solution algorithm in a tighter space, leading to a shorter runtime. After meeting the convergence criteria or reaching the maximum number of iterations, we output the final values of $(\mathbf{p}^{*},\mathbf{r}^{*},\mathbf{x}^{*},\mathbf{v}^{*})$.
@@
1:Input: Initial upper-level variable $\mathbf{p}^{0},\mathbf{r}^{0}$, initial penalty parameter $\rho^{0}$, bound relaxation coefficient $\zeta$, penalty parameter update coefficient $\psi^{-},\psi^{+}$, optimality gap target $\tau$, objective tolerance $\epsilon$, time limit $T$, maximum iterations $K$
@@
4:  Step 1: For given $\mathbf{p}^{k},\mathbf{r}^{k}$, solve the lower-level problem Eq. (1) to obtain optimal $\mathbf{x}^{k},\mathbf{v}^{k}$